# CoARA will not save science from the tyranny of administrative evaluation

### v2

@@ 4 Philip Kitcher’s well-ordered science
The introduction of the ideal discussion allows Kitcher to individuate two emerging dangers in defining the relevant issues to be investigated. According to Kitcher, in the well-ordered science of a democratic society, no one should have the authority to decide what the relevant issues are (for a Kantian perspective see Pievatolo [2024]). The relevant issues cannot be identified by a simple majority voting procedure, because this could lead to the emergence of the “tyranny of ignorance”, where myopic voters choose “in ignorance of the possibilities, and of the consequences for others, completely absorbed in their own self-directed wishes". The second danger is the “tyranny of unwarranted expertise” that occurs if researchers are given the right not only to do research, but also to choose which topics should be investigated. Well-ordered science takes care to avoid these two tyrannies in the definitions of relevant issues to be investigated.
@@ 5 Administrative evaluation of research and well-ordered science
To answer the question, one can start from the observation that in administrative procedures of research evaluation, evaluators are academics with administrative authority. Consider, for example, the evaluation of a research article in a massive research assessment. This evaluation is neither the result of a well-ordered certification nor the result of a public discussion among peers. It is instead the indisputable judgment of one or more scientists invested of the administrative authority to act as evaluators. As highlighted by Pievatolo [2024], the fact that evaluators are scientists does not signify that they are ‘peer’ of the scientists being evaluated, since only the former have administrative authority. The evaluated scientists can challenge the judgment they receive not through reasoning in public debate but only through administrative or judicial action. In Italy, for example, to challenge the judgment received in the national research assessment (VQR) or in the procedures for the qualification (ASN), a researcher must resort to a court action before an administrative court. In short, evaluators act as ’guardians of science’ for the policy makers who have granted them the authority to evaluate.
@@
Let us now imagine, alternatively, that the government authority chooses evaluators from among academics, that judgments are based on “originality, significance and rigour”, and that the final score of each article is computed by more or less complex procedures (average of evaluations, majority vote or other). In this case, paraphrasing Kitcher, it is not difficult to think of the emergence of the tyranny of unwarranted expertise: A particular ideological agenda, a pervasive bias, pushes members of a scientific subcommunity vested with the authority to evaluate to overestimate the quality and impact of certain products and to underestimate others even while fabricating the necessary evidence. For instance, evidence exist showing that the tyranny of unwarranted expertise was at work in British [Lee and Harley, 1998, Lee et al., 2013] and Italian [Baccini et al., 2020, Baccini and De Nicolao, 2021, Baccini and Re, 2024] research assessments.
@@
6 Conclusive remarks
Administrative evaluation of research is not a component of well-ordered science, it is not a certification, but an activity that attaches value to research products according to the criteria decided by administrative authorities and applied by scientists who are turned into administrative officials. By its mere implementation, administrative evaluation of research distances science from Kitcher’s ideal of well-ordered science in a democratic society.
CoARA is focused on evaluation technology. With DORA and the Leiden Manifesto, it shares the idea that research evaluation is needed, and that the adoption of the ‘good’ evaluative technology permits to achieve socially desirable outcomes. The same idea is also shared by scholars who advocate the opposite view, that the best technology is the adoption of well-designed bibliometric indicators [Abramo, 2024, for all]. The defense of informed peer review as the best technology and that of bibliometric indicators are but two sides of the same coin. The discussion on the best technology for administrative evaluation of research is firmly founded in the neo-liberal vision of how science should be organized and managed. The alternative proposals do not consider a relevant problem for contemporary science the authoritarian character of administrative evaluation of research. While proponents of bibliometrics usually do not hide the technocratic nature of their vision, CoARA instead hides the authoritarian nature of the evaluation behind the screen of peer review supported by a responsible use of metrics.
@@ 6 Conclusive remarks
CoARA’s commitments accept to place the authority of evaluation in the hands of a few academics capable of performing peer review supported by a responsible use of metrics. Similarly, advocates of the use of bibliometrics suggest placing the authority of evaluation in the hands of a few academics capable of handling bibliometrics [Abramo, 2024]. From this perspective, the “forced battle between peer review and scientometric research assessment” [Abramo, 2024] does not appear as an intellectual battle. It appears, instead, as a battle for the authority to evaluate, to occupy positions in government offices or agencies that conduct administrative evaluation of research, and to sell well-paid evaluative services. In fact, the growing space between policy makers and scientific system is progressively filled by academics transformed in professionals of evaluation, by consultancy agencies and data providers [Jappe, 2020]. The winners in the battle for evaluation technology will have access to the market for the expertise they claim to possess, in the same vein as the illiberal reformers told by Leonard [2016].
@@
The obvious objection to the reasoning so far is that Kitcher’s ideal conversation is an impractical idealization, since the organization of contemporary science institutions requires administrative evaluation. The substitution of technology proposed by CoARA and the recommendations contained in the Leiden Manifesto could therefore be considered as realistic and feasible proposals to reform administrative evaluation of research by shrinking or at least by correcting the use of metrics.
@@
The counter-objection starts from the observation that emergence of pervasive administrative evaluation of research is relatively recent and is the result of the extension of the recipes of new public management in the realm of science policy [Sandström and Van den Besselaar, 2018]. Science systems had worked for centuries without the new public management recipes and without the administrative evaluation of research, being organized differently in different countries [Gillies, 2008, Sandström and Van den Besselaar, 2018, Whitley, 2007]. Hence, a possible option for research policy may be to abandon the precepts of new public management and to return to limit the administrative evaluation of research to a bare minimum. This minimum could consist in limiting the administrative evaluation of research to only those procedures strictly necessary for the basic functioning of science, i.e. the recruitment of researchers and the funding of project-based research. This means eliminating performance-based funding mechanisms for research institutions, massive individual and aggregate evaluation procedures, and direct cash incentives for publications and citations. The main advantage of this elimination is to greatly reduce not only the risks of tyranny of expertise associated with, but also the distortions in scientists’ behavior induced by administrative evaluation of research.
Even if the administrative evaluation of research is strictly limited, the simple adoption of peer review, as proposed by CoARA, is not sufficient to completely eliminate the risk of tyranny of expertise. When administrative evaluation of research is used for the recruitment of researchers and the funding of projects, the main problem is the design of a fair procedures [Beersma and De Dreu, 2003, Baccini and Re, 2024], minimizing the risk of tyranny of expertise.
Procedural fairness is focused on the social system and procedures which generate evaluation results, and not on the final distribution of rewards or punishments. This approach is not consequentialist: the social desirability of a procedure is based on its fairness, and not in its consequences, such as the distribution of rewards or the effects on research quality and impact. Designing the evaluation procedure in a way that respects the principle of fairness is essential if the results of the evaluation and the distribution of rewards, whatever they may be, are to be fair, or at least not unfair. Fairness regards the consistency of the procedure over time and between persons, “its accuracy and prevention of personal bias; or its representativeness of the values, interests, and outlook of important subgroups in the population of persons affected by the allocative process” [Leventhal, 1980].
Administrative evaluation of research, even when limited as proposed above, and conducted by peer review, as proposed by CoARA, attributes the power to decide to a panel, i.e. to a specific group of scholars. The most critical issues, to reduce to a minimum tyranny of expertise, is respecting fairness in the composition of the evaluation panel and limit its decision power. For example, when selecting experts for a panel, it is necessary to consider not only easily observable dimensions such as the gender or affiliation of panel members, but also the plurality of research approaches that exist in a given field [Baccini and Re, 2024]. In the case of the distribution of resources for research projects, a fair procedure can be designed by incorporating some random elements for reducing biases induced by decision power of panelists [Avin, 2018]. For instance, Gillies [2008], among the others, suggests the adoption of a two-stage process in which experts select a set of fundable projects from which those to be funded are drawn at random.
The limitation of the domain in which administrative evaluation of research is applied, and the design of fair procedures may contribute to reduce the risks of the tyranny of ignorance and the tyranny of expertise in contemporary systems of science. A goal that CoARA and, more generally, any proposal focusing only on evaluation technologies cannot achieve.

### v3

@@ 4 Philip Kitcher’s well-ordered science
The introduction of the ideal discussion allows Kitcher to individuate two emerging dangers in defining the relevant issues to be investigated. According to Kitcher, in the well-ordered science of a democratic society, no one should have the authority to decide what the relevant issues are (for a Kantian perspective see Pievatolo [2024a]). The relevant issues cannot be identified by a simple majority voting procedure, because this could lead to the emergence of the “tyranny of ignorance”, where myopic voters choose “in ignorance of the possibilities, and of the consequences for others, completely absorbed in their own self-directed wishes". The second danger is the “tyranny of unwarranted expertise” that occurs if researchers are given the right not only to do research, but also to choose which topics should be investigated. Well-ordered science takes care to avoid these two tyrannies in the definitions of relevant issues to be investigated.
@@ 5 Administrative evaluation of research and well-ordered science
To answer the question, one can start from the observation that in administrative procedures of research evaluation, evaluators are academics with administrative authority. Consider, for example, the evaluation of a research article in a massive research assessment. This evaluation is neither the result of a well-ordered certification nor the result of a public discussion among peers. It is instead the indisputable judgment of one or more scientists invested of the administrative authority to act as evaluators. As highlighted by Pievatolo [2024a], the fact that evaluators are scientists does not signify that they are ‘peer’ of the scientists being evaluated, since only the former have administrative authority. The evaluated scientists can challenge the judgment they receive not through reasoning in public debate but only through administrative or judicial action. In Italy, for example, to challenge the judgment received in the national research assessment (VQR) or in the procedures for the qualification (ASN), a researcher must resort to a court action before an administrative court. In short, evaluators act as ’guardians of science’ for the policy makers who have granted them the authority to evaluate.
@@
Let us now imagine, alternatively, that the government authority chooses evaluators from among academics, that judgments are based on “originality, significance and rigour”, and that the final score of each article is computed by more or less complex procedures (average of evaluations, majority vote or other). In this case, paraphrasing Kitcher, it is not difficult to think of the emergence of the tyranny of unwarranted expertise: A particular ideological agenda, a pervasive bias, pushes members of a scientific subcommunity vested with the authority to evaluate to overestimate the quality and impact of certain products and to underestimate others even while fabricating the necessary evidence. For instance, evidence exist showing that the tyranny of unwarranted expertise was at work in British [Lee and Harley, 1998, Lee et al., 2013] and Italian [Baccini et al., 2020, Baccini and De Nicolao, 2021, Baccini and Re, 2025] research assessments.
@@
6 Discussion and conclusion
The application of Kitcher’s concept of well-ordered science in the context of a democratic society allows for a clearer understanding of administrative research evaluation as an activity fundamentally external to the domain of science itself. Rather than serving as a scientific certification process, administrative evaluation functions as a mechanism for ascribing value to research outputs and contributions based on criteria established by administrative or policy authorities. These criteria are then implemented by scientists who, in this context, assume the role of administrative agents, rather than autonomous members of a scientific community. As such, the very act of implementing administrative evaluation shifts the practice of science away from Kitcher’s ideal of a well-ordered science, in which research is shaped by an ideal collective deliberation informed by both epistemic values and democratic input. In contrast, administrative evaluation imposes externally-defined priorities and bureaucratic standards, subject both to tyranny of ignorance and to tyranny of expertise.
The research assessment reform movement, and CoARA in particular, originates from the premise that administrative research evaluation is both necessary and inevitable. In the case of CoARA, this foundational assumption is reflected not only in its stated objectives but also in its institutional genesis. As discussed above, despite the appearance of broad consultation and stakeholder engagement, the initiative was largely orchestrated by the European Commission, with key orientations, procedures, and design features defined in advance by a limited group of actors. This continuity is not incidental. Since the Lisbon Strategy of 2000, when the European Union declared its ambition to become “the most competitive and dynamic knowledge-based economy in the world”, a succession of policy frameworks, including Europe 2020 and the renewed European Research Area (ERA), have institutionalized administrative research evaluation as a central governance mechanism. Now that these strategies appear to have difficulties to deliver on their transformative promises, CoARA can be understood as an effort to re-legitimize and adapt the evaluative apparatus to new discourses around openness, transparency, and inclusivity, without relinquishing its underlying logic.
The core problem CoARA seeks to address lies in the overreliance on quantitative metrics, such as citation counts and journal impact factors, which has led to widespread goal displacement: researchers optimize for what is measured rather than for what truly matters to scientific progress or societal benefit. The reform agenda is grounded in the belief that by adopting more appropriate, inclusive, and context-sensitive evaluative technologies is possible to realign research incentives and foster outcomes that are both scientifically meaningful and socially desirable. According to CoARA, the most promising technology consists of peer-review informed by the responsible use of metrics, that can be readily adapted to the diverse contexts and purposes of different evaluation procedures.
In turn, the notion that administrative research evaluation is necessary is also supported by scholars who advocate for the use of carefully constructed bibliometric indicators [Abramo, 2024, for all]. In this perspective, the defense of informed peer review and that of bibliometrics are, in effect, two sides of the same coin. Both aim to enhance the administrative research evaluation, albeit through different means. This debate about alternative evaluative technologies is deeply embedded in a neo-liberal conception of science governance. Within this vision, research is organized and managed as a productive system whose outputs must be continuously monitored, evaluated, and optimized for “maximise the quality and impact of research” [Coalition for Advancing Research Assessment (CoARA), 2022, p. 1].
The two alternative proposals fail to recognize the authoritarian nature of administrative research evaluation as a structural problem within contemporary science. At its core, this model concentrates evaluative authority in the hands of a limited group of scholars, who are thereby empowered to “put an end to the free exchange of ideas and override the legitimate opinions” of other scientists [Kitcher, 2011]. CoARA’s commitments explicitly endorse the delegation of evaluative authority to selected academics deemed capable of conducting peer review, supported by the “responsible” use of metrics. Similarly, proponents of bibliometrics advocate for entrusting evaluation to expert bibliometricians [Abramo, 2024]. In this latter case, the technocratic character of research assessment is not only acknowledged but openly defended: evaluation must be governed by experts in quantitative indicators. By contrast, CoARA obscures the authoritarian dimension of evaluation by framing it within the language of peer review, cloaked in ideals of collegiality and responsibility. Yet, in both cases, evaluative power remains exclusionary exposing research assessment to the risks of a tyranny of expertise, the removal of legitimacy from dissenting scientific perspectives and reinforcement of hierarchical control over what counts as valuable research.
@@ 6 Discussion and conclusion
From this perspective, the “forced battle between peer review and scientometric research assessment” [Abramo, 2024] does not appear to be an intellectual battle. Rather, it appears as a battle for the authority to evaluate, to occupy positions in government offices or agencies that conduct administrative evaluation of research, and to sell well-paid evaluative services. In fact, the growing space between policy makers and scientific system is progressively filled by academics transformed in professionals of evaluation, by consultancy agencies and data providers [Jappe, 2020]. The winners n this battle over evaluation technologies will gain privileged access to the market for expertise they claim to possess, mirroring the trajectory of the illiberal reformers described by Leonard [2016]. They turned technocratic authority into a platform for political and economic power. Leonard shows how, in the early 20th century, a new class of experts, economists, administrators, and policy advisors, used the language of science and rationality to displace democratic deliberation with technocratic control, often serving elite interests under the guise of neutrality.
@@
The obvious objection to the reasoning so far is that Kitcher’s ideal conversation is only an impractical idealization, since the organization of contemporary science institutions necessarily requires administrative evaluation. The substitution of technology proposed by CoARA and the recommendations contained in the Leiden Manifesto could therefore be considered as a realistic and feasible proposals to reform administrative evaluation of research by shrinking or at least by correcting the use of metrics.
The counter-objection begins by noting that the emergence of pervasive administrative research evaluation is a relatively recent phenomenon [Sandström and Van den Besselaar, 2018]. Until the 1980s, national science systems operated without New Public Management mechanisms and without systematic administrative evaluation, relying instead on diverse organizational models shaped by specific historical and institutional contexts [Gillies, 2008, Gingras, 2016, Sandström and Van den Besselaar, 2018, Whitley, 2007]. The exponential growth of contemporary science does not, in itself, justify the necessity of these administrative tools. As many critics have argued [Gillies, 2008, Collini, 2012], the problems that gave rise to the research assessment reform movement point not so much to the need for better evaluative technologies, but to the need for a more fundamental rethinking of the purpose, utility, and necessity of administrative research evaluation itself.
@@
A possible option for research policy may be to abandon the precepts of new public management and to return to limit the administrative evaluation of research to a bare minimum. This minimum could consist in limiting the administrative evaluation of research to only those procedures strictly necessary for the basic functioning of science: the recruitment of researchers and the funding of project-based research. This means eliminating performance-based funding mechanisms for research institutions, massive individual and aggregate evaluation procedures, and direct cash incentives for publications and citations. The main advantage of this elimination is to greatly reduce not only the risks of tyranny of expertise and ignorance associated with, but also the distortions in scientists’ behavior induced by administrative evaluation of research.
Even in a scenario where administrative research evaluation is strictly limited, the use of peer review, as proposed by CoARA, does not fully eliminate the risk of a tyranny of expertise. When such evaluation is employed for purposes like researcher recruitment or project funding, the central challenge becomes the design of fair procedures [Beersma and De Dreu, 2003, Baccini and Re, 2025], aimed at minimizing the concentration of evaluative power and reducing the risk of epistemic exclusion. Procedural fairness emphasizes the integrity of the evaluation process itself, independently of its outcomes. It is a non-consequentialist notion: what makes a procedure fair is not the distribution of rewards it produces or its effects on research quality, but its internal characteristics, such as consistency, impartiality, and the inclusion of diverse perspectives and interests of those affected by the evaluation [Leventhal, 1980].
Even when administrative research evaluation is narrowly applied, as suggested here, and implemented through peer review, as proposed by CoARA, it still concentrates decision-making power in the hands of a panel, that is, a specific group of scholars. To minimize the risks of a tyranny of expertise, two conditions are crucial: ensuring fairness in the composition of evaluation panels and carefully limiting their authority [Beersma and De Dreu, 2003, Baccini and Re, 2025]. For instance, selecting panel members should not rely solely on visible dimensions like gender or institutional affiliation, but must also reflect the plurality of epistemic approaches within a field [Baccini and Re, 2025]. In the allocation of research funding, fairness can also be enhanced by incorporating elements of randomization to mitigate bias and reduce the discretionary power of reviewers [Avin, 2018]. Gillies [2008] and others, for instance, have suggested, a two-stage process, where experts identify a pool of meritorious projects, from which final selections are made at random.
Ultimately, limiting the scope of administrative research evaluation and designing fair evaluative procedures can help mitigate both the tyranny of expertise and the tyranny of ignorance in contemporary science systems. This is a goal that CoARA, and, more broadly, any reform agenda focused exclusively on improving evaluative technologies, cannot achieve on its own, unless it is accompanied by a more fundamental rethinking of the role, purpose, and governance of research evaluation itself.